You are graduating from college at the end of this semester and have decided to invest $5,000 at the end of each year into a Roth IRA for the next 30 years. If you earn 6% compounded annually on your investment of $5,000 at the end of each year, how much will you have when you retire in 30 years? How much will you have if you wait 10 years before beginning to save and only make 20 payments into your retirement account?
Answer #1

PMT =5000
Number of years =30
Rate =6%
FV using annuity formula=PMT*(((1+r)^n-1)/r)=5000*(((1+6%)^30-1)/6%)=395,290.93

If Number of years =20
FV using annuity formula=PMT*(((1+r)^n-1)/r)=5000*(((1+6%)^20-1)/6%)=183927.96